OpenStudy (anonymous):

A $200.00 sweater has been marked down to $120.00. What percentage discount from the original price does the markdown represent?

OpenStudy (anonymous):

120 / 200 = .6 then .6 * 100% = 60% They are paying 60% so you are saving 40%

OpenStudy (anonymous):

40% because the markdown is $80/$200

OpenStudy (anonymous):

40%

OpenStudy (anonymous):

\[\frac{200-120}{200}*100 =40\% \] AngelPSP and orn2440 get an A